# A Midsummer Meme’s Dream: Investigating Market Manipulations in the Meme Coin Ecosystem

### v1

@@ 3.1.3. Dataset Refinement
Manual inspection of high market cap or high-priced tokens in our dataset reveals some false positives, such as stablecoins or wrapped tokens. We apply a three-stage filtering process to eliminate misclassified tokens and retain only meme coins, minimizing false positives even at the cost of excluding some legitimate ones. This conservative approach prioritizes precision over recall, enhancing the reliability of our subsequent analyses.
@@
First, we use CoinMarketCap’s API (CoinMarketCap, 2025a) to identify and remove stablecoins. By cross-referencing our dataset against CoinMarketCap’s list of 82 labeled stablecoins, we discard 4 tokens, reducing our dataset to 35,116 tokens.
@@
In the second refinement stage, we focus on identifying non-meme tokens with significant market presence. Specifically, we examine tokens that meet either of two conditions: a price above \$0.80—since meme coins typically have very low prices (Nani, 2022), this threshold helps capture stablecoins that may be slightly depegged—or a market capitalization exceeding $10^{7}$, as the cumulative distribution function (CDF) in Fig. 8 (Appendix) shows that most tokens in our dataset fall below this value. This process identifies 525 tokens for detailed review. We then use CoinGecko’s API (CoinGecko, 2025a) to retrieve their associated categories and remove those labeled with non-meme classifications (e.g., Staking, Bridged). This step results in the removal of 46 additional tokens.
@@
Finally, we apply a conservative string-matching filter to remove any remaining tokens with names containing terms typically associated with non-meme assets: specifically usd, wrapped, and staked. This step identifies and excludes 182 tokens, resulting in a final dataset of 34,988 meme coins. Tab. 1 provides an overview of the tokens collected from various sources and the filtering steps applied to refine the final dataset.
Table 1. Summary of meme coin data collection.
@@
[…]
This table summarizes the meme coins data collection and filtering process
3.2. Dataset Overview
We collected the data in mid-October 2024, capturing a comprehensive snapshot of the meme coin ecosystem. For each of the 34,988 tokens in our dataset, we gather technical details, including the token name, symbol, contract address, and blockchain. In particular, the contract address and blockchain serve as unique identifiers, enabling the tracking of tokens across different platforms.
3.2.1. Blockchain Distribution
@@
Ethereum (Ethereum.org, 2025) was the first platform to support smart contracts, laying the foundation for decentralized applications. However, its high fees and scalability limitations led to the development of alternatives. One was the creation of new Layer 1 blockchains like BNB Smart Chain (BSC) (Bnbchain.org, 2024), which is EVM compatible and offers lower transaction costs. Another was the rise of Layer 2 solutions built on Ethereum itself, such as Base (Base.org, 2025), which improves scalability while remaining part of the Ethereum ecosystem. In contrast, Solana (Foundation, 2025b) is not EVM compatible and takes a different approach, prioritizing low fees and high transaction throughput through its unique proof-of-history consensus mechanism. Together, these four blockchains account for 32,876 meme coins, representing 93.96% of our dataset. Given their dominance, our subsequent analysis focuses on tokens deployed on these platforms.
3.2.2. Addresses Validation
As the next step, we perform a validation process to identify and discard invalid addresses and non-deployed tokens (e.g., placeholders such as 0x000comingsoon or upcoming). For EVM-compatible blockchains (BSC, Ethereum, and Base), we first validate address format: they must be 42 characters long, start with ”0x,” and contain only hexadecimal characters. This step identifies 791 invalid addresses (739 in BSC, 51 in Ethereum, 1 in Base). Next, for addresses with the correct format, we use the explorers’ APIs (Etherscan (Etherscan, 2025) for Ethereum, BscScan (BscScan, 2025) for BSC, and BaseScan (BaseScan, 2025) for Base) to check for their creation transactions. We find 254 addresses, 211 in BSC and 43 in Ethereum, that lacked creation transactions and, thereby, are not deployed contracts. For Solana tokens, we use Solana’s RPC API (Foundation, 2025a) to verify valid mint addresses, identifying 20 invalid addresses.
@@ 3.2.2. Addresses Validation
Table 2. Summary of token address validation process
Blockchain Invalid Not Deployed Removed Remaining
@@
3.2.3. Temporal Evolution of Meme Coin Creation.
To analyze the creation patterns of the meme coins over time, we collect their contract deployment timestamps. For EVM-compatible blockchains, we use their explorers’ APIs (Etherscan, BscSca, and BaseScan). Concerning Solana, we leverage the API offered by Alchemy (Alchemy Insights, 2025), a blockchain development service that provides tools for building decentralized applications.
@@
Late 2020 Bull Run. The cryptocurrency bull market, marked by Bitcoin’s rise from \$10,000 to over \$60,000 (Martin Young, 2021), coincided with two major developments: the launch of SHIBA INU and the introduction of Binance Smart Chain (BSC), that significantly boosted meme coin creation thanks to its low transaction fees (Coin98, 2023).
Pepe Coin & BONK. In April 2023, the launch of Pepe coin on Ethereum challenged BSC’s dominance. This was followed by the surge of BONK in October 2023, driven by the introduction of single-sided staking pools (Compass, 2024).
Pump.fun. The peak in meme coin creation came in the summer of 2024, especially in June and July, when Solana emerged as the leading platform. This surge is largely due to the rapid rise in popularity of pump.fun (Boaz Sobrado, 2025).
Figure 1. Creation date of meme coins over time and their distribution on different blockchains.
[…]
A plot showing the creation date of meme coins over time and their distribution on different blockchains.
4. Tokenomics and Price Growth
@@
4.1. Tokenomic Anatomy Across Blockchains
[…]
A figure containing three CDF plots comparing meme coins across Ethereum, BSC, Solana, and Base. The first plot shows the distribution of price, the second shows the distribution of circulating supplies, and the third shows the distribution of market cap.
@@ 4.1. Tokenomic Anatomy Across Blockchains
(b) Circulating Supply
(c) Market Cap
Figure 2. Distribution of the prices (2(a)), circulating supplies (2(b)), and market cap (2(c)) of the meme coins on Ethereum, BSC, Solana, and Base grouped according to their creation date.
@@
Of the 31,811 meme coins collected, we successfully obtained economic information for 20,692 meme coins (65.05%) in US dollars. The remaining 11,119 tokens for which we cannot retrieve additional data have been collected on DexScreener, Coinsniper, and CoinGecko. These platforms derive token price, circulating supply, and market capitalization from liquidity pools where tokens are traded. However, if a token lacks a liquidity pool at the time of data collection, its price and market cap cannot be determined. This explains why no such information is available for these 11,119 tokens. After removing four tokens with a price of zero (all on BSC), the final dataset consists of 20,688 meme coins: 10,468 on Solana, 7,493 on BSC, 2,107 on Ethereum, and 620 on Base.
@@
We group the meme coins in our dataset according to when they were created to examine how their age correlates with their economic performance and characteristics. Specifically, we group tokens into quarterly periods for 2024 ($Q1_{2024}$, $Q2_{2024}$, $Q3_{2024}$, and $Q4_{2024}$), with an additional category for tokens created before 2024 (pre-2024). Fig. 2 illustrates how economic metrics vary across these periods for each blockchain platform.
4.1.1. Price Structures
@@
The price analysis reveals an ecosystem characterized by tokens with very low prices, alongside a few high-priced outliers. The overall distribution shows a median price of \$$4\times 10^{-6}$ with approximately 85% of tokens priced below \$0.00001. This pricing strategy creates a psychological illusion of accessibility: investors can acquire billions of tokens for minimal capital, fostering the narrative that these assets could replicate the trajectories of Dogecoin or Shiba Inu (Nani, 2022). Moreover, we identify 14 meme coins with prices exceeding \$100, each employing one of two distinct price strategies. The first involves artificial scarcity mechanisms using regular token burning, as seen in tokens like Ponzio The Cat (valued at approximately \$20 billion) and HayCoin (valued at \$250,000). These tokens have been flagged as scams by security platforms such as Token Sniffer (Sniffer, 2025) and show systematic supply reduction designed to inflate prices, rewarding early investors at the expense of those who enter later. The second approach involves the use of single-token supply models, as seen in tokens such as Potato (valued at \$55,000). This extreme scarcity is leveraged to justify their unusually high valuations, despite limited utility or demand.
Concerning the temporal analysis of the price, Fig. 2(a)) reveals three distinct blockchain-specific pricing patterns.
Higher Prices for Newer Tokens on Ethereum and BSC: These platforms exhibit sequential price increases across time periods, suggesting a market preference for newly launched tokens. Pre-2024 tokens show median prices of $6\times 10^{-8}$ (Ethereum) and $5\times 10^{-10}$ (BSC), while $Q4_{2024}$ tokens reach $4\times 10^{-5}$ (Ethereum) and $1\times 10^{-4}$ (BSC), representing three to six orders of magnitude increase.
Cyclical pricing on Base: Tokens on Base demonstrate non-monotonic price progression, starting with relatively high values in early periods ($2\times 10^{-5}$ for pre-2024), declining through mid-2024 (to $2\times 10^{-6}$ in $Q3_{2024}$), and recovering to $1\times 10^{-5}$ by $Q4_{2024}$.
Stable pricing on Solana: Conversely, Solana demonstrates remarkable price consistency across all creation periods, maintaining a median of approximately $8\times 10^{-6}$ regardless of token age.
4.1.2. Token Supply Strategies
An important element of the tokenomics of a crypto token is its circulating supply, the total number of tokens currently available on the market (not locked or burned). While established cryptocurrencies like Bitcoin (21 million coins) and Ethereum (about 120 million) emphasize supply scarcity, meme coins often implement hyperinflated supply models: 43.29% of meme coins in our dataset present circulating supplies exceeding 1 billion tokens, with 30.68% exceeding 100 billion tokens. Even at the 25th percentile, meme coin supplies reach 973 million tokens. This extreme supply inflation serves a psychological purpose: it allows promoters to highlight the appeal of owning billions of tokens, distracting from their actual low value. As shown in Fig. 2(b), supply strategies vary significantly across blockchains.
Hyperinflated supplies on Ethereum and BSC: These platforms host tokens presenting high variability with median supplies often exceeding $10^{11}$.
Moderate supplies on Base: Base tokens typically maintain supplies in the $10^{9}$-$10^{11}$ range, with significant but less extreme variation compared to Ethereum and BSC.
Standardized supplies on Solana: Solana exhibits consistent token supplies concentrated around $10^{9}$, due to the standardized creation processes implemented by pump.fun, which applies fixed supply parameters.
4.1.3. Market Capitalization
@@
Market capitalization, calculated by multiplying the circulating supply of a token by its current price, is a key indicator of the market value of a project. Among the meme coins we analyzed, the median market cap is just \$4,100, highlighting their highly speculative nature. Interestingly, a few meme coin projects report extremely high market capitalizations. While some well-known tokens, such as Dogecoin (\$53 billion), Shiba Inu (\$15 billion), and Pepe (\$8 billion), have substantial and legitimate market value, these are rare exceptions in a space often marked by value manipulation. For example, catgirl a claims an implausible market cap of \$151 billion despite having almost no liquidity, suggesting how unreliable market capitalization can be as a standalone metric for meme coins. Upon investigating the 47 meme coins reporting market caps above \$1 billion, we detect significant evidence of market manipulation. More than half have liquidity under \$1,000, and 88.1% show liquidity below 0.1% of their reported market cap. This discrepancy points to a common tactic: inflating token prices through small, strategic trades in low-liquidity environments, a form of manipulation detailed in Sec. 5.3.
Regarding the cross-chain analysis, Fig. 2(c) suggests consistent temporal patterns in market cap across all blockchains: newer tokens (those created in $Q4_{2024}$) present dramatically higher market caps compared to older ones. Specifically, Base shows the highest median market capitalization for recent tokens at \$882,507, followed by Ethereum at \$443,688 and Solana at \$214,981. BSC tokens remain significantly lower at \$12,223, approximately one order of magnitude below their counterparts on other blockchains.
@@
This strong preference for newly launched tokens is more consistent across blockchains than pricing patterns, suggesting that token novelty may be a stronger driver of investor capital allocation than the underlying blockchain infrastructure.
4.1.4. Cross-Platform Economic Patterns
@@
4.2. Three-Month Price Performance Evaluation
To study the price evolution of meme coins, we tracked their prices over three months following our initial data collection until mid-January 2025.
4.2.1. Data Availability and Retrieval
@@
Of the 361 tokens with missing price data, we originally collected these from multiple sources: 12 from Gecko Terminal, 350 from DexScreener, and 21 from Coingecko (with some overlap between platforms). As discussed in Sec. 4.1.1, platforms such as Gecko Terminal and DexScreener require tokens to maintain active liquidity pools to determine prices. Thus, the absence of price data for these tokens likely indicates their liquidity pools were removed during our study period. Meanwhile, Coingecko may delist tokens that fail to meet its listing criteria (CoinGecko, 2025d), particularly those with poor liquidity or absent development teams. The blockchain with the highest percentage of tokens lacking updated price data is Ethereum with 185 tokens (8.78%), followed by BSC with 138 (1.84%), Base with 8 (1.29%), and Solana with 30 (0.29%).
4.2.2. Price Change Distribution
Fig. 3 presents the cumulative distribution functions (CDFs) of meme coin price changes across different blockchains. For clarity, tokens with returns exceeding 100% are capped at 100% in the visualization. Specifically, 6,938 tokens (34.13%) experience price decreases, 6,462 tokens (31.79%) maintain stable prices, and 6,927 tokens (34.07%) see price increases.
Figure 3. CDFs of meme coin price changes by blockchain platform over three months.
[…]
A plot showing the CDFs of the price changes of meme coins grouped by different blockchains (Ethereum, BSC, Solana, and Base).

### v2

@@ 3.1.3 Dataset Refinement
Manual inspection of high market cap or high-priced tokens reveals some false positives, such as stablecoins or wrapped tokens. We apply a three-stage filtering process to eliminate misclassified tokens and retain only meme coins, minimizing false positives even at the cost of excluding some legitimate ones. This conservative approach prioritizes precision over recall, enhancing the reliability of our subsequent analyses.
@@
First, we use CoinMarketCap’s API [coinmakretcap_api] to identify and remove stablecoins. By cross-referencing our dataset against CoinMarketCap’s list of 82 labeled stablecoins, we discard 4 tokens, reducing our dataset to 35,116 tokens.
@@
In the second refinement stage, we focus on identifying non-meme tokens with significant market presence. Specifically, we examine tokens that meet either of two conditions: a price above \$0.80—since meme coins typically have very low prices [nani2022doge], this threshold helps capture stablecoins that may be slightly depegged—or a market capitalization exceeding $10^{7}$, as the cumulative distribution function (CDF) in Fig. 8 (Appendix) shows that most tokens in our dataset fall below this value. This process identifies 525 tokens for detailed review. We then use CoinGecko’s API [coingecko_api] to retrieve their associated categories and remove those labeled with non-meme classifications (e.g., Staking, Bridged). This step results in the removal of 46 additional tokens.
@@
Finally, we apply a conservative string-matching filter to remove tokens with names containing terms typically associated with non-meme assets: specifically usd, wrapped, and staked. This step identifies and excludes 182 tokens, resulting in a final dataset of 34,988 meme coins, which we publicly release [meme_coin_repository]. Tab. 1 provides an overview of the tokens collected and the filtering steps applied to refine the dataset.
Table 1: Summary of meme coin data collection.
@@
3.2 Dataset Overview
We collected the data from all targeted aggregators in mid-October 2024. For each of the 34,988 tokens in our dataset, we store the token name, symbol, contract address, and blockchain. In particular, the contract address and blockchain serve as unique identifiers, enabling the tracking of tokens across different platforms.
Concerning data coverage and limitations, we note that tokens sourced from CoinMarketCap and CoinGecko include aggregated data from both Centralized (CEX) and Decentralized Exchanges (DEX). For tokens identified solely through DEX aggregators, our view is limited to DEX. Nevertheless, this does not introduce coverage gaps: If a meme coin is sufficiently established to be listed on a CEX, it is likely indexed by aggregators such as CoinMarketCap. Consequently, tokens found exclusively on DEX aggregators are early-stage assets that have not yet been listed on centralized platforms.
3.2.1 Blockchain Distribution
@@
Ethereum [ethereum_doc] was the first platform to support smart contracts, laying the foundation for decentralized applications. However, its high fees and scalability limitations led to the development of alternatives. One was the creation of new Layer 1 blockchains like BNB Smart Chain (BSC) [bnbsmartchain], which is EVM compatible and offers lower transaction costs. Another was the rise of Layer 2 solutions built on Ethereum itself, such as Base [base_doc], which improves scalability while remaining part of the Ethereum ecosystem. In contrast, Solana [solana_doc] is not EVM compatible and takes a different approach, prioritizing low fees and high transaction throughput through its proof-of-history consensus mechanism. Together, these four blockchains account for 32,876 meme coins, representing 93.96% of our dataset. Given their dominance, our subsequent analysis focuses on tokens deployed on these platforms.
3.2.2 Addresses Validation
As the next step, we perform a validation process to identify and discard invalid addresses and non-deployed tokens (e.g., placeholders such as 0x000comingsoon or upcoming). For EVM-compatible blockchains (BSC, Ethereum, and Base), we first validate address format: they must be 42 characters long, start with "0x," and contain only hexadecimal characters. This step identifies 791 invalid addresses (739 in BSC, 51 in Ethereum, 1 in Base). Next, for addresses with the correct format, we use the explorers’ APIs (Etherscan [etherscan_api] for Ethereum, BscScan [bscscan_api] for BSC, and BaseScan [base_scan] for Base) to check for their creation transactions. We find 254 addresses, 211 in BSC and 43 in Ethereum, that lacked creation transactions and, thereby, are not deployed contracts. For Solana tokens, we use Solana’s RPC API [solana_rpc] to verify valid mint addresses, identifying 20 invalid addresses.
@@ 3.2.2 Addresses Validation
Table 2: Summary of token address validation process
Platform Invalid Undeployed Removed Retained
@@
3.2.3 Temporal Evolution of Meme Coin Creation.
To analyze the creation patterns of the meme coins, we collect their contract deployment timestamps. For EVM-compatible blockchains, we use their explorers’ APIs (Etherscan, BscSca, and BaseScan). Concerning Solana, we leverage the API offered by Alchemy [alchemy_api], a blockchain development service that provides tools for building decentralized applications.
@@
Late 2020 Bull Run. The cryptocurrency bull market, marked by Bitcoin’s rise to over \$60,000 [bull_run_2021], coincided with two major events: the launch of SHIBA INU and the introduction of Binance Smart Chain (BSC), that significantly boosted meme coin creation thanks to its low transaction fees [bsc_success].
Pepe Coin & BONK. In April 2023, the launch of Pepe coin on Ethereum challenged BSC’s dominance. This was followed by the surge of BONK in October 2023, driven by the introduction of single-sided staking pools [bonk_staking_pool].
Pump.fun. The peak in meme coin creation came in the summer of 2024, especially in June and July, when Solana emerged as the leading platform. This surge is largely due to the rapid rise in popularity of pump.fun [pump_success].
Figure 1: Creation date of meme coins over time and their distribution on different blockchains.
4 Tokenomics and Price Growth
@@
4.1 Tokenomic Anatomy Across Blockchains
@@
(b) Market Cap
Figure 2: Distribution of the prices (2(a)) and market cap (2(b)) of the meme coins. Ethereum and BSC show sequential growth, with recent tokens ($Q4_{2024}$) reaching valuations orders of magnitude higher than pre-2024 baselines. Solana maintains stable pricing but shares the trend of significantly higher market caps for new launches, while Base exhibits high variability without clear temporal patterns.
@@
Of the 31,811 meme coins collected, we successfully obtained economic information for 20,692 meme coins (65.05%) in US dollars during our data snapshot in mid-October 2024. The remaining 11,119 tokens for which we cannot retrieve additional data have been collected on DexScreener, Coinsniper, and CoinGecko. These platforms derive token price, circulating supply, and market capitalization from liquidity pools where tokens are traded. However, if a token lacks a liquidity pool at the time of data collection, its price and market cap cannot be determined. This explains why no such information is available for these 11,119 tokens. After removing four tokens with a price of zero (all on BSC), the final dataset consists of 20,688 meme coins: 10,468 on Solana, 7,493 on BSC, 2,107 on Ethereum, and 620 on Base.
@@
We group the meme coins in our dataset according to when they were created to examine how their age correlates with their economic characteristics. Specifically, we group tokens into quarterly periods for 2024 ($Q1_{2024}$, $Q2_{2024}$, $Q3_{2024}$, and $Q4_{2024}$), with an additional category for tokens created before 2024 (pre-2024). To validate the observed market behaviors, we performed a non-parametric Kruskal-Wallis H test [kruskal1952use], followed by Dunn’s post-hoc analysis with Bonferroni correction. We consider differences to be statistically significant only when $p<0.01$ [dunn1961multiple]. Fig. 2 illustrates how economic metrics vary across these periods for each blockchain.
4.1.1 Price Structures
@@
The price analysis reveals an ecosystem characterized by tokens with very low prices, alongside a few high-priced outliers. The overall distribution shows a median price of $4\times 10^{-6}$ with approximately 85% of tokens priced below \$0.00001. This pricing strategy creates a psychological illusion of accessibility: investors can acquire billions of tokens for minimal capital, fostering the narrative that these assets could replicate the trajectories of Dogecoin or Shiba Inu [nani2022doge].
Cases of Suspiciously High Pricing: We identify 14 meme coins with prices exceeding \$100, each employing one of two distinct price strategies. The first involves artificial scarcity mechanisms using regular token burning, as seen in tokens like Ponzio The Cat (valued at approximately \$20 billion) and HayCoin (\$250,000). These tokens have been flagged as scams by security platforms such as Token Sniffer [token_sniffer] and show systematic supply reduction designed to inflate prices, rewarding early investors at the expense of those who enter later. The second approach involves the use of single-token supply models, as seen in tokens such as Potato (valued at \$55,000). This extreme scarcity is leveraged to justify their unusually high valuations, despite limited utility.
Concerning the temporal analysis of the price, Fig. 2(a)) reveals that Ethereum and BSC exhibit sequential price increases across time periods, suggesting a market preference for newly launched tokens. Base shows high price variability without statistically significant temporal patterns. Conversely, Solana demonstrates price consistency across all creation periods, maintaining a median of approximately $8\times 10^{-6}$ regardless of token age.
4.1.2 Market Capitalization
Market capitalization, calculated by multiplying the circulating supply of a token by its current price, is a key indicator of the market value of a project. Among the meme coins we analyzed, the median market cap is just \$4,100, highlighting their low market cap and highly speculative nature.
@@ 4.1.2 Market Capitalization
Cases of High Market Caps with Low Liquidity: A few meme coin projects report extremely high market cap. While some well-known tokens, such as Dogecoin (\$53 billion), Shiba Inu (\$15 billion), and Pepe (\$8 billion), have substantial and legitimate market value, these are rare exceptions in a space often marked by value manipulation. For example, catgirl a claims an implausible market cap of \$151 billion despite having almost no liquidity, suggesting how unreliable market capitalization can be as a standalone metric for meme coins. Upon investigating the 47 meme coins reporting market caps above \$1 billion, we detect significant evidence of market manipulation. More than half have liquidity under \$1,000, and 88.1% show liquidity below 0.1% of their reported market cap. This discrepancy points to a common tactic: inflating token prices through small, strategic trades in low-liquidity environments, a form of manipulation detailed in Sec. 5.3.
@@
Regarding the cross-chain analysis (Fig. 2(b)), we found Solana and Ethereum show statistically significant increases in market cap for $Q4_{2}024$ tokens, with Ethereum tokens reaching a median of \$443K and Solana tokens \$214K. BSC shows a similar but weaker trend with $Q4_{2}024$ tokens at \$12K, while Base exhibits high variability without consistent temporal patterns. This strong preference for newly launched tokens is more consistent across blockchains than pricing patterns, suggesting that token novelty may be a stronger driver for investors than the underlying blockchain infrastructure.
4.1.3 Cross-Platform Economic Patterns
@@
4.2 Price Performance Evaluation
To study the price evolution of meme coins over three-months following our initial snapshot, we conducted a follow-up data collection in mid-January 2025.
4.2.1 Data Availability and Retrieval
@@
Concerning the 361 tokens with missing price, we originally collected them from multiple sources (with some overlap): 12 from Gecko Terminal, 350 from DexScreener, and 21 from Coingecko. As discussed in Sec. 4.1.1, Gecko Terminal and DexScreener require tokens to maintain active liquidity pools to determine prices. Thus, the absence of price data for these tokens likely indicates their liquidity pools were removed during our study period. Meanwhile, Coingecko may delist tokens failing to meet its listing criteria [coingecko_listing], particularly those with poor liquidity. The blockchain with the highest percentage of tokens lacking updated price is Ethereum with 185 tokens (8.78%), followed by BSC with 138 (1.84%), Base with 8 (1.29%), and Solana with 30 (0.29%).
4.2.2 Price Change Distribution
Fig. 3 presents the cumulative distribution functions (CDFs) of meme coin price performance across different blockchains. To visualize the extreme disparity between losses and gains, we use a logarithmic scale for the price multiplier, defined as $Price_{final}/Price_{initial}$. Specifically, 6,938 tokens (34.13%) experience price decreases, 6,462 tokens (31.79%) maintain stable prices, and 6,927 tokens (34.07%) see price increases.
Figure 3: Log-CDFs of meme coin price changes by blockchain over three months. Many tokens experience almost no price change, suggesting they are likely inactive.